OpenStudy (anonymous):

Find the standard form of the equation of the line that passes through (1/2,-2) and has slope -4.

jigglypuff314 (jigglypuff314):

standard form of a line: ax + by = c first you want to put what you know into a equation you know either point-slope form: y - y1 = m(x - x1) or slope-intercept form: y = mx + b

OpenStudy (anonymous):

would it be point-slope? @jigglypuff314

jigglypuff314 (jigglypuff314):

if you know point-slope, plug in and tell me what you get because once you get that then you would solve so that the values with x and y are alone on one side.

jigglypuff314 (jigglypuff314):

point-slope form: y - y1 = m(x - x1) given (1/2,-2) and has slope -4 y + 2 = (-4)(x - (1/2)) distribute y + 2 = (-4)(x) - (-4)(1/2) simplify y + 2 = -4x + 2 subtract 2 from both sides -2 -2 y = -4x add 4x to both sides 4x + y = 0 so that would be your answer If you don't understand something, feel free to ask :)

OpenStudy (anonymous):

how would u know when to use point slope and slope intercept form? @jigglypuff314

jigglypuff314 (jigglypuff314):

it would be a preference or what your teacher had taught you

OpenStudy (anonymous):

OH okay thank you